calcuate the mass needed of Al to create 130 kg of Cr from Cr2O3


1
Expert's answer
2021-04-21T06:41:17-0400

Cr2O3 + 2Al = 2Cr + Al2O3


M(Cr) = 52 kg/kmol

M(Al) = 27

kg/kmol


n(Cr2O3) = 130 kg / 52 kg/kmol

= 2.5kg/mol

2n(Al) = 2n(Cr) =

2.5*27 = 67.5

= 67.5Kg